prove the following :                 1) sin0(1+tan0)+cos0(1+cot0)=sec0+cosec0.   please give me some tips amd suggestions so that i can prove tough problems and please say how to practice trigonomety and when ?

Grade:10

2 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
sin\theta (1+tan\theta ) + cos\theta (1+cot\theta ) = sec\theta +cosec\theta
sin\theta (1+tan\theta ) + cos\theta (1+cot\theta ) = sin\theta (1+tan\theta ) + \frac{cos^{2}\theta}{sin\theta } (1+tan\theta )(1+tan\theta )(\frac{cos^{2}\theta +sin^{2}\theta }{sin\theta }) = \frac{1+tan\theta }{sin\theta } = sec\theta+cosec\theta
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty
Crappy
5 years ago
Shitty the day of the day of the day of school prayer also have to take its course of action and avoid using the same was true that was last modified the morning's or